OpenStudy (anonymous):

Multiply five-eighths times twenty. thirty-two twelve and one-half 2 and one-half one thirty-second

OpenStudy (anonymous):

B. Twelve and One-Half. 5*20 = 100/8 = 12 1/2

OpenStudy (anonymous):

can u help me with more @*Shadow*

OpenStudy (anonymous):

I got 12 and a half to

OpenStudy (camerondoherty):